Grade 12th passPhysical Chemistry

For the reaction: 2HI(g) = H2(g) + I2(g)
The degree of dissociation(a) of HI(g) is related to equilibrium constant, Kp by the expression
  1. { 2 (2Kp)1÷2 } ÷ {1+ 2 (Kp)1÷2 }
  2. {1 + 2 (Kp)1÷2 } ÷ 2
  3. { (1 + 2Kp) ÷ 2 }1÷2
  4. { 2Kp ÷ (1 + 2Kp) }1÷2
Please please give answer step by step
NOTE : PLEASE give step by step answer...

Yasir Rather
7 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
2HI             →         H2         +            I2
t=teq.            C(1-a)                         Ca/2                          Ca/2
Kp = (Ca/2)^2/(C(1-a))^2 => a^2/4(1-a)^2
a/(1-a) = 2root(Kp)
a = 2root(Kp) – 2root(Kp)*a
a(1 + 2root(Kp)) = 2root(Kp)
a = 2root(Kp)/(1 + 2root(Kp))
 
Regards
Arun (askIITians forum expert)